OpenStudy (anonymous):

Write an equation in slope-intercept form for a line that passes through the two points (-2, 7) and (-4, -3). Show your work.

OpenStudy (jack1):

slope-intercept form : y = mx +c now m = gradient and c = y-intercept

OpenStudy (jack1):

so @res2712 , do u know how to find gradient from 2 points?

OpenStudy (anonymous):

No I don't .

OpenStudy (jack1):

ok, so gradient (m) is rise / run \[\large m = \frac {y_2 - y_1}{x_2 - x_1}\]

OpenStudy (jack1):

so u have 2 points: \[\large (-2, 7)~ and~ (-4, -3)\] which are:\[\large (x_1, y_1)~ and~ (x_2, y_2)\]. so if \[\large m = \frac {y_2 - y_1}{x_2 - x_1}\]... m = ?

OpenStudy (anonymous):

-2-(-4) 7-(-3) ?

OpenStudy (jack1):

yeah... but reverse that... so \[\large m = \frac {y_2 - y_1}{x_2 - x_1}\] \[\large m = \frac {-3 - 7}{-4 - (-2)}\] =???

OpenStudy (anonymous):

-10/-6?

OpenStudy (jack1):

almost, its a double minus on the bottom row, so = -10/-2 = 10/2 = 5

OpenStudy (anonymous):

So that's my answer?

OpenStudy (jack1):

for m so equation is y = mx + c to solve question, u need values for m and c

OpenStudy (jack1):

y = mx + c and m = 5, so y = 5x + c now, do u know how to find C?

OpenStudy (anonymous):

I really don't